QUANTITATIVE IR SPECTRA AND VIBRATIONAL ASSIGNMENTS OF CH_{2}I_{2}, AN ATMOSPHERIC AEROSOL PRECURSOR

T.J. JOHNSON, T. MASIELLO, S.W. SHARPE, Pacific Northwest National Laboratory, P.O. Box 999, MSIN K8-88, Richland, WA 99354.

As part of the Northwest Infrared (NWIR) database of quantitative infrared spectra, we have recently completed quantitative spectra of diiodomethane, CH_{2}I_{2}. Photolysis of this molecule in the presence of ozone, O_{3}, has been suggested as an immediate precursor to new particle formation over the oceans, particularly in coastal areas. Combined with the quantitative medium (0.1 cm^{−1}) resolution vapor-phase IR spectra, liquid-phase IR and FT-Raman spectra, as well as ab initio calculations have been used to update and extend the vibrational assignments of earlier workers. Two strong b_{2} symmetry A-type bands at 584 and 1114 cm^{−1} are observed, but are not resolved at 760 Torr and appear as B-type. In contrast, the b_{1} symmetry C-type bands near 5953, 4426 and 3073 cm^{−1} are resolved with rotational structure, including Q-branches with widths < 1 cm^{−1}. Potential use of these bands for atmospheric monitoring will be discussed.